OpenStudy (anonymous):

Find the first term of the following geometric sequence _____, 42, 126, 378, 1,134, ...

OpenStudy (geneticrockhopper247):

The trick here is to find the pattern. Since it is a geometric sequence, divide the third number by the second to find the multiple.

OpenStudy (anonymous):

so is the anser 3

OpenStudy (geneticrockhopper247):

42*3=126. 126*3=378 378*3=1134. Correct :)

OpenStudy (geneticrockhopper247):

Do you understand???

OpenStudy (anonymous):

yea thanks

OpenStudy (geneticrockhopper247):

you're welcome